OpenStudy (anonymous):

the length is 3 more inches than the width the perimeter is 30 inches

OpenStudy (anonymous):

length is 9 inches and the width is 6 inches lol did that in my head

OpenStudy (kkutie7):

so I'm guessing this is a rectangle? =) if so P=l+l+w+w or P=2l+2w l=3+w P=30

OpenStudy (kkutie7):

30=2l+2w 30=2(3+w)+2w 30=6+2w+2w 30=6+4w 24=4w 6=w

OpenStudy (kkutie7):

l=3+w l=3+6 l=9